OpenStudy (anonymous):

Check my answer? Identifying compression or shrink? Identify a horizonal or vertical stretch or compression of the function f(x) = 1/x by observing the equation of the function g(x)=5/x (A)A horizontal stretch by a factor of 5. (B)A vertical stretch by a factor of 5. (C)A horizontal compression by a factor of 5. (D)A vertical compression by a factor of 5. I believe the answer is C because a division shrinks and the x axis lies horizontally.

ganeshie8 (ganeshie8):

g(x) = 5/x = 5*(1/x) = 5*f(x)

ganeshie8 (ganeshie8):

thats clearly a vertical stretch by a factor of 5 it could also be a horizontal compression by a factor of 1/5 but this is not in your options

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

thank you...sucks my answer was the complete opposite!

ganeshie8 (ganeshie8):

http://www.mathsisfun.com/sets/function-transformations.html this site has nice explanation

ganeshie8 (ganeshie8):

you need to remember these by heart while working with transformations : http://gyazo.com/e21d11e1d6df67afe8bd4c656d98260a

OpenStudy (anonymous):

Can you also tell me how to identify odd/even/neither functions @ganeshie8 ?

OpenStudy (anonymous):

Pretty sure this is an even function though